OpenStudy (anonymous):

Find the angle between the given vectors to the nearest tenth of a degree. u = <-5, -4>, v = <-4, -3> i got -9.1 am i right?

OpenStudy (dumbcow):

did you use \[\cos \theta = \frac{u*v}{|u| |v|}\]

OpenStudy (anonymous):

no? @dumbcow

OpenStudy (dumbcow):

ok well that is the equation you need to use for finding angle between vectors

OpenStudy (dumbcow):

do you know how to find dot product and magnitude?

OpenStudy (anonymous):

not really. @dumbcow